Theere are 156 blocks to make a wall. The bottom 3 rows will all have the same number of blocks. The next 6 rows will each have 2 fewer than the row before it. How many blocks are in each row? I know its 22 for the bottom 3 and I know the rest just not the way to write equation
Answer by addingup(3677)

Let's call the number of blocks in each of the first three rows x. Then we have:
3x, since the first three rows each have x blocks.
After that, the rows will have x-2 followed by x-2-2= x-4 and so on, like this:
3x+x-2+x-4+x-6+x-8+x-10+x-12=156
9x-42=156
9x=156+42
9x=198
x=198/9
x=22 is the number of blocks in the first 3 rows.
then you'll have:
20,18,16,14,12 & 10 in each successive row.
proof
3*22+20+18+16+14+12+10=156
66+90=156
156=156 We have the correct answer.